Solve for X and Y in the following problems using either substitution or elimination methods. Make sure you show all your work so you can get partial credit even if your final answer is wrong.
a. X + Y= 6 , 2X + Y = 8
b. 7X + 3Y = 14 , 5X + 9Y = 10
c. 4X + Y = 16 , 2X + 3Y = 24
d. 12X + Y = 25 , 8X - 2Y = 14

Answer by checkley77(12844)

a.
X + Y = 6
-(2X + Y = 8) SUBTRACT.
------------
-X=-2
X=2 ANS.
-2+Y=6
Y=6+2
Y=8 ANS.
2*2+Y=8
4+Y=8
Y=8-4
Y=4 ANS.
PROOF:
2+4=6
6=6
2*2+4-8
4+4=8
8=8
b.
7X + 3Y = 14 MULTIPLY BUY -3 & ADD
5X + 9Y = 10
-21X-9Y=-42
------------------
-16X=-32
X=-32/-16
X=2 ANS.
7*2+3Y=14
14+3Y=14
3Y=14-14
3Y=0
Y=0 ANS.
PROOF:
5*2+9*0=10
10=10
c.
4X + Y = 16 MULTIPLY BY -3 & ADD
2X + 3Y = 24
-12X-3Y=-48
-----------------
-10X=-24
X=-24/-10
X=2.4 ANS.
4*2.4+Y=16
9.6+Y=16
Y=16-9.6
Y=6.4 ANS.
PROOF:
2*2.4+3*6.4=24
4.8+19.2=24
24=24
d.
12X + Y = 25 MULTIPLY BY 2 & ADD .
8X - 2Y = 14
24X+2Y=50
---------------
32X=64
X=64/32
X=2 ANS.
12*2+Y=25
24+Y=25
Y=25-14
Y=1 ANS.
PROOF:
8*2-2*1=14
16-2=14
14=14